OpenStudy (anonymous):

-√72 explain step by step

OpenStudy (anonymous):

72 faactors into 36*2 and 36 is a perfect square, so when you apply the root to it a 6 comes outside the radical\[-\sqrt{72}=-\sqrt{36*2}=-\sqrt{36}*\sqrt{2}=-6\sqrt{2}\]